# On endosplit $p$-permutation resolutions and Broué’s conjecture for $p$-solvable groups

### v2

@@ 4. Galois descent of endosplit $p$-permutation resolutions
This follows since both every component $\hat{C}_{i}$ or differential $d_{i}:\hat{C}_{i}\to\hat{C}_{i-1}$ has some finite field $l$ that realizes $\hat{C}_{i}$ or $d_{i}$, since $k^{\prime}$ is the colimit of the diagram $\mathbb{F}_{p}\hookrightarrow\mathbb{F}_{p^{2}}\hookrightarrow\dots$, so any matrix representation or intertwiner will have values contained in some finite field. Therefore, there exists a finite field $l$ which realizes $\hat{C}$, i.e. there exists a chain complex $C^{\prime}$ of $lG$-modules for which $k^{\prime}\otimes_{l}C^{\prime}\cong\hat{C}$. By Theorem 4.8, $C^{\prime}$ descends to $k$, as desired. ∎
@@ 5. Constructing splendid Rickard equivalences for blocks of $p$-solvable groups over $\mathbb{F}_{p}$
In this section, we apply the results of the previous section to show that Kessar and Linckelmann’s refinement of Broué’s abelian defect group conjecture holds for blocks of $p$-solvable groups $G$ satisfying $G=O_{p^{\prime},p,p^{\prime}}(G)$ with abelian defect groups.
@@
[KL18] “Given a categorical equivalence, say a Morita equivalence or a Rickard equivalence between $\mathcal{O}^{\prime}Gb$ and $\mathcal{O}Hc$ for some complete discrete valuation ring $\mathcal{O}^{\prime}$ for $G$ and $H$ finite groups, $b$ and $c$ blocks of $\mathcal{O}^{\prime}G$ and $\mathcal{O}^{\prime}H$ respectively, and a complete discrete valuation ring $\mathcal{O}$ contained in $\mathcal{O}$, is the equivalence between $\mathcal{O}^{\prime}Gb$ and $\mathcal{O}^{\prime}Hc$ an extension of an equivalence between $\mathcal{O}Gb$ and $\mathcal{O}Hc$? … [This] may be viewed as evidence for a refined version of the Abelian defect group conjecture, namely that for any $p$-modular system $(K,\mathcal{O},k)$ and any block $b$ of $\mathcal{O}G$ with abelian defect group $P$ and Brauer correspondent $c$, there is a splendid Rickard equivalence between $\mathcal{O}Gb$ and $\mathcal{O}N_{G}(H)c$.”
We do so by showing that Harris and Linckelmann’s construction in [HL00] can be adapted to a $p$-modular system with $k=\mathbb{F}_{p}$. One may use unique lifting of splendid Rickard equivalences to lift to $\mathcal{O}=\mathbb{Z}_{p}$, which then implies the result for any extension of the $p$-modular system $(\mathbb{Q}_{p},\mathbb{Z}_{p},\mathbb{F}_{p})$. In [HL00], the authors specify a $p$-modular system with $k$ algebraically closed. However, for many of the statements, one may replace the assumption that $k$ is algebraically closed with the assumption that $k$ is perfect with enough roots of unity, which we will use frequently.
For this section, to avoid confusion, we will refer to “blocks” as the primitive idempotents of $Z(\overline{\mathbb{F}}_{p}G)$ (note these always exist over some finite field), and refer to block idempotents over a finite field $k$ which may not be primitive upon extending scalars as “block idempotents of $kG$.” Note that for these block idempotents, defect groups are defined in the usual way (for instance, a maximal $p$-subgroup $P$ of $G$ for which $\operatorname{Br}_{P}(b)\neq 0$) and satisfy essentially all the usual properties. Brauer’s first main theorem asserts that there is a canonical bijection between the set of block idempotents of $kG$ with defect group $P$ and the set of block idempotents of $kN_{G}(P)$ with defect group $P$. Given a block idempotent $b$ of $kG$ with defect group $P$, the block idempotent $c$ of $kN_{G}(P)$ corresponding to $b$ under this bijection is called the Brauer correspondent.
@@
An $\mathcal{O}^{\times}$-group is a group $\hat{G}$ endowed with a group homomorphism $\mathcal{O}^{\times}\to\hat{G}$ whose image lies in $Z(\hat{G})$; we usually denote by $\hat{\lambda}$ the image of $\lambda\in\mathcal{O}^{\times}$ in $\hat{G}$, and by $\check{G}$ the opposite $\mathcal{O}^{\times}$-group, which as an abstract group is equal to $\hat{G}$ but endowed with the homomorphism $\mathcal{O}^{\times}\to\check{G}$ sending $\lambda\in\mathcal{O}^{\times}$ to $\check{\lambda}=(\hat{\lambda})^{-1}.$
@@
For our context, $\mathcal{O}^{\times}$-groups arise from the action of $G$ on a matrix algebra $S$ over $\mathcal{O}$ by algebra automorphisms. Denote by ${}^{x}s$ the action of $x\in G$ on $s\in S$. By the Skolem-Noether theorem, any automorphism of $S$ is inner, and hence for any $x\in G$, there is $s\in S$ such that ${}^{x}t=sts^{-1}$ for all $t\in S$. The group $\hat{G}$ of all such pairs $(x,s)\in G\times S^{\times}$ satisfying ${}^{x}t=sts^{-1}$ for all $t\in S$ for all $t\in S$ becomes an $\mathcal{O}^{\times}$-group with the group homomorphism sending $\lambda\in\mathcal{O}^{\times}$ to $\hat{\lambda}=(1_{G},\lambda 1_{S})$. We call $\hat{G}$ the $\mathcal{O}^{\times}$-group defined by the action of $G$ in $S$. In fact, $\hat{G}$ is a central $\mathcal{O}^{\times}$-extension of $G$, as we have a short exact sequence
$1\to\mathcal{O}^{\times}\to\hat{G}\to G\to 1$
@@
[Lin18a, Corollary 5.3.4] Let $(K,\mathcal{O},k)$ be a $p$-modular system with $k$ algebraically closed. If $S$ has $\mathcal{O}$-rank prime to $p$ and $P$ is a finite $p$-group acting on a matrix algebra $S$, the action of $P$ lifts to a unique group homomorphism $\sigma:P\to S^{\times}$ such that $\det(\sigma(x))=1$ for all $x\in P$. In particular, $\hat{P}$ splits canonically as a direct product $\hat{P}\cong P\times\mathcal{O}^{\times}$, with the map sending $u\in P$ to $(u,\sigma(u))\in\hat{P}$ being a section of the canonical surjection $\hat{P}\to P$.
@@
For every $S$ and $P$ satisfying the hypotheses in the previous proposition, there exists a $p$-modular system $(K,\mathcal{O},k)$ with $k$ finite for which the previous proposition holds as well. See [Lin18a, Remark 5.3.11] and [Lin18a, Remark 1.2.19] for more details.
@@
If $\hat{G}$ is an $\mathcal{O}^{\times}$-group, we denote by $\mathcal{O}_{*}\hat{G}$ the quotient of the group algebra $\mathcal{O}\hat{G}$ by the ideal generated by the set of elements $\lambda.\hat{x}-1_{\mathcal{O}}.(\hat{\lambda}\hat{x})$ where $\lambda$ runs over $\mathcal{O}$ and $\hat{x}$ runs over $\hat{G}$. $\mathcal{O}_{*}\hat{G}$ is $\mathcal{O}$-free of rank $|G|$, since the image in $\mathcal{O}_{\ast}\hat{G}$ of any subset of order $|G|$ of $\hat{G}$ mapping onto $G$ under the canonical surjection $\hat{G}\to G$ is an $\mathcal{O}$-basis of $\mathcal{O}_{*}\hat{G}$.
Let $(K,\mathcal{O},k)\subseteq(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be an extension of $p$-modular systems, let $S$ be a matrix algebra over $\mathcal{O}$, and set $S^{\prime}:=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}S$. Let $\hat{G}_{\mathcal{O}}$ (resp. $\hat{G}_{\mathcal{O}^{\prime}}$) denote the $\mathcal{O}$-group (resp. $\mathcal{O}^{\prime}$-group) defined by the action of $G$ on $S$ (resp. $G$ on $S^{\prime}$). Then, the obvious inclusion of $\mathcal{O}^{\prime}$-algebras $\mathcal{O}^{\prime}\otimes_{\mathcal{O}}\mathcal{O}_{*}\hat{G}_{\mathcal{O}}\to\mathcal{O}_{*}^{\prime}\hat{G}_{\mathcal{O}^{\prime}}$ is an isomorphism, since it sends the $\mathcal{O}^{\prime}$-basis given by representatives of $G$ in $\hat{G}$ to the analogous $\mathcal{O}^{\prime}$-basis.
@@
[HL00, Statements 2.5, 2.6] If $\hat{G}$ is a central $\mathcal{O}^{\times}$-extension of a finite group $G$, there is a unique natural algebra homomorphism $\mathcal{O}G\to\mathcal{O}_{\ast}\hat{G}\otimes_{\mathcal{O}}\mathcal{O}_{\ast}\check{G}$ mapping $x\in G$ to $\hat{x}\otimes\hat{x}$, where $\hat{x}$ is any element of $\hat{G}$ (and hence also of $\check{C}$) which lifts $x$. Denote by $\operatorname{Ind}^{\hat{G}\times\check{G}}_{\Delta G}:{}_{\mathcal{O}G}\mathbf{mod}\to{}_{\mathcal{O}_{\ast}\hat{G}\otimes_{\mathcal{O}}\mathcal{O}_{\ast}\check{G}}\mathbf{mod}$ the obvious induction functor. Note any $\mathcal{O}_{\ast}\hat{G}\otimes_{\mathcal{O}}\mathcal{O}_{\ast}\check{G}$-module is equivalently a $(\mathcal{O}_{\ast}\hat{G},\mathcal{O}_{\ast}\hat{G})$-bimodule since $(\mathcal{O}_{\ast}\check{G})^{o}\cong\mathcal{O}_{\ast}\hat{G}$.
Furthermore, for any two $\mathcal{O}G$-modules $U,V$ there is a unique natural isomorphism of $(\mathcal{O}_{*}\hat{G},\mathcal{O}_{*}\hat{G})$-bimodules
@@
To prove Broué’s conjecture holds, Harris and Linckelmann provide a reduction to the case where $G=O_{p^{\prime},p,p^{\prime}}(G)$, then use structural properties of groups of this form to construct a splendid Rickard autoequivalence built from a Morita equivalence and a diagonally induced endosplit $p$-permutation resolution. We will verify that their construction for $G=O_{p^{\prime},p,p^{\prime}}(G)$ descends to the block’s field of realization. We remark that in [HL00], the construction is performed over $\mathcal{O}$, however, we will work in positive characteristic instead, then invoke Rickard’s lifting theorem, [Ric96, Theorem 7.2].
We first restate the structural theorem that describes block decompositions in the case of $G=O_{p^{\prime},p,p^{\prime}}(G)$. In the original paper, this is stated over $\mathcal{O}$, but it is straightforward that this also holds over $k$.
@@
    The action of $P$ on $S$ lifts to a unique group homomorphism $\sigma:P\to S^{\times}$ such that $\det(\sigma(x))=1$ for all $x\in P$, and there is a $p^{\prime}$-subgroup $E$ of $\operatorname{Aut}(P)$ and a central $k^{\times}$-extension $\hat{E}$ of $E$ such that we have an isomorphism of $k^{\times}$-groups
@@
The assumption that $k$ is algebraically closed is invoked for Proposition 5.4, [HL00, Lemma 2.7] which is a consequence of Dade’s splitting theorem of fusion in endopermutation modules, and in the assumption that $S$ is a matrix algebra. One may verify Dade’s theorem satisfies the same large enough finite field replacement condition via a similar argument; see the proof of [Lin18b, Theorem 7.9.1] and [Lin18a, Remark 1.2.19]. Therefore, there exists a $p$-modular system $(K,\mathcal{O},k)$ with $k$ finite (large enough for $G$) for which the previous proposition holds as well.
@@
Therefore, we have a splendid Morita equivalence between $kHc$ and $k_{\ast}\hat{L}$, since source algebra equivalences induce splendid Morita equivalences (see [Lin01, Theorem 4.1], note that the notion of splendid-ness used in this article is stronger than Rickard’s original notion). Therefore, to verify Broué’s conjecture, it suffices to construct a splendid Rickard equivalence for $kGb$ and $k_{\ast}\hat{L}$. To verify the refined abelian defect group conjecture, it suffices to show that the equivalence descends to $\mathbb{F}_{p}[b]=\mathbb{F}_{p}[c]$ (see [BY22, Lemma 4.2]), then apply [KL18, Theorem 6.5(b)].
@@
    $D_{k}(G)=D^{\Omega}(G)$, where $D^{\Omega}(G)$ is the subgroup of $D_{k}(G)$ generated by equivalence classes of relative syzygies, that is, kernels of augmentation homomorphisms $kX\to k$, where $X$ is a transitive $G$-set.
@@
    If $M$ is an indecomposable endopermutation $kG$-module, then $M$ is absolutely indecomposable.
In particular, every endopermutation $kP$-module descends to $\mathbb{F}_{p}$ (or any intermediate field extension), that is, there exists an endopermutation $\mathbb{F}_{p}G$-module $\tilde{M}$ such that $k\otimes_{\mathbb{F}_{p}}\tilde{M}\cong M$.
@@
Statement (a) follows from the classification in [Bou06, Theorem 9.5] (expanding on the original classification for abelian $p$-groups in [Dad78]), and statement (b) follows from [Dad78, Theorem 6.6]. For the final statement, see [Lin18b, Remark 7.8.5]. ∎
@@
We now discuss the construction of the splendid Rickard equivalence between $kGb$ and $kHc$ and prove descent. We continue to use the notation of Proposition 5.8. As noted in Remark 5.9, it suffices to construct a splendid Rickard equivalence between $kGb\cong kO_{p^{\prime}}(G)b\otimes_{k}k_{\ast}\hat{L}$ and $k_{\ast}\hat{L}$. There are two components of this construction, an $S$-module $V$ (which as $kP$-module is endopermutation) and a Rickard tilting complex (non-splendid) of $(k_{\ast}\hat{L},k_{\ast}\hat{L})$-bimodules $X$. We describe the construction of both of them, and show that they both descend to $\mathbb{F}_{p}[b]$.
@@
    By Lemma 5.10, the $kP$-module $V$ descends to $l$ (in fact, it descends to $\mathbb{F}_{p}$). Let $\tilde{V}$ denote the endopermutation $lP$-module for which $k\otimes_{l}\tilde{V}\cong V$. Then the structural homomorphism $P\to\operatorname{End}_{k}(V)^{\times}$ has image contained in the $l$-subalgebra $\operatorname{End}_{l}(\tilde{V})$, therefore under the interior $P$-algebra isomorphism $S\cong\operatorname{End}_{k}(V)$, the image of $\sigma:P\to S^{\times}$ is contained in the $l$-subalgebra $\tilde{S}=lO_{p^{\prime}}(G)b$. In particular, we have an isomorphism of interior $P$-algebras $S\cong k\otimes_{l}\tilde{S}$. Therefore, we have an isomorphism of interior $P$-algebras $\tilde{S}\cong\operatorname{End}_{l}(\tilde{V})$. It follows again that $\tilde{V}$ is an $\tilde{S}$-module and that we have $\operatorname{End}_{l}(\tilde{V})\cong\tilde{S}$ as $(\tilde{S},\tilde{S})$-bimodules. Thus, the $S$-module $V$ descends to $l$ via $\tilde{V}$.
  • Since $V$ is an endopermutation $kP$-module with $P$ abelian, by [Ric96, Theorem 7.2] $V$ has an endosplit $p$-permutation resolution $X_{V}$. The uniqueness of the lifting $\sigma:P\to S^{\times}$ of the action of $P$ on $S$ with determinant 1 implies the isomorphism class of $V$ is $E$-stable, and thus $X_{V}$ is an $E$-stable endosplit $p$-permutation resolution. Then $V^{*}\otimes_{k}X_{V}$ is an endosplit (non-$p$-permutation) resolution of $V^{*}\otimes_{k}V$, and it follows from [HL00, Lemma 1.5(i)] that $\operatorname{Ind}^{L}_{P}(V^{*}\otimes_{k}X_{V})$ is an endosplit resolution of $\operatorname{Ind}^{L}_{P}(V^{*}\otimes_{k}V).$ Since $V$ is capped, $V^{*}\otimes_{k}V$ has the trivial $kP$-module as a direct summand, thus $\operatorname{Ind}^{L}_{P}(V^{*}\otimes_{k}V)$ has the trivial $kL$-module as a direct summand, since $E$ is a $p^{\prime}$-group. Therefore, $\operatorname{Ind}^{L}_{P}(V^{*}\otimes_{k}X_{V})$ has a direct summand $Y$ which is an endosplit resolution of the trivial $kL$-module. Setting
@@
    and this is a split complex with homology concentrated in degree 0 isomorphic to $\operatorname{Ind}^{\hat{L}\times\check{L}}_{\Delta L}(k)\cong k_{\ast}\hat{L}$. Thus $X$ is a Rickard tilting complex of $(k_{\ast}\hat{L},k_{\ast}\hat{L})$-bimodules.
    Since $V$ descends to $\mathbb{F}_{p}$, $X_{V}$ descends to $l$ by Theorem 4.8 (note that it remains $E$-stable since Theorem 4.2 states that descent is unique). By construction, $Y$ also descends to $l$. $k_{\ast}\hat{L}$ also descends to $l$ since $k\otimes_{l}l_{\ast}\hat{L}\cong k_{\ast}\hat{L}$, so the functor $\operatorname{Ind}^{\hat{L}\times\check{L}}_{\Delta L}$ is well defined, hence $X$ also descends to a chain complex $\tilde{X}$ of $(l_{\ast}\hat{L},l_{\ast}\hat{L})$-bimodules satisfying $\tilde{X}^{*}\otimes_{l_{\ast}\hat{L}}\tilde{X}\simeq l_{\ast}\hat{L}$.
It follows that $V\otimes_{k}X$ is a Rickard tilting complex of $(S\otimes_{k}k_{\ast}\hat{L},k_{\ast}\hat{L})$-bimodules with homology concentrated in degree 0 isomorphic to $V\otimes_{k}k_{\ast}\hat{L}$. Then $V\otimes_{k}X$ is verified to be splendid on [HL00, Page 95], completing the proof for the case $G=O_{p^{\prime},p,p^{\prime}}(G)$. We have shown above that $V\otimes_{k}X$ descends to a Rickard tilting complex $\tilde{V}\otimes_{l}\tilde{X}$ of $(\tilde{S}\otimes_{l}l_{\ast}\hat{L},l_{\ast}\hat{L})$-bimodules with homology concentrated in degree 0 isomorphic to $\tilde{V}\otimes_{l}l_{\ast}\hat{L}$. The same proof also shows that $\tilde{X}\otimes_{l}\tilde{V}$ is splendid. From this we obtain the following partial result.

### v3

@@ 4. Galois descent of endosplit $p$-permutation resolutions
This follows since both every component $\hat{C}_{i}$ or differential $d_{i}\colon\hat{C}_{i}\to\hat{C}_{i-1}$ has some finite field $l$ that realizes $\hat{C}_{i}$ or $d_{i}$, since $k^{\prime}$ is the colimit of the diagram $\mathbb{F}_{p}\hookrightarrow\mathbb{F}_{p^{2}}\hookrightarrow\dots$, so any matrix representation or intertwiner will have values contained in some finite field. Therefore, there exists a finite field $l$ which realizes $\hat{C}$, i.e. there exists a chain complex $C^{\prime}$ of $lG$-modules for which $k^{\prime}\otimes_{l}C^{\prime}\cong\hat{C}$. By Theorem 4.8, $C^{\prime}$ descends to $k$, as desired. ∎
@@ 5. Constructing splendid Rickard equivalences for blocks of $p$-solvable groups over $\mathbb{F}_{p}$
In this section, we apply the results of the previous section to show that Kessar–Linckelmann’s refinement of Broué’s abelian defect group conjecture holds for certain blocks of $p$-solvable groups $G$ satisfying $G=O_{p^{\prime},p,p^{\prime}}(G)$ with abelian defect groups.
@@
[KL18] “Given a categorical equivalence, say a Morita equivalence or a Rickard equivalence between $\mathcal{O}^{\prime}Gb$ and $\mathcal{O}^{\prime}Hc$ for some complete discrete valuation ring $\mathcal{O}^{\prime}$ for $G$ and $H$ finite groups, $b$ and $c$ blocks of $\mathcal{O}^{\prime}G$ and $\mathcal{O}^{\prime}H$ respectively, and a complete discrete valuation ring $\mathcal{O}$ contained in $\mathcal{O}^{\prime}$, is the equivalence between $\mathcal{O}^{\prime}Gb$ and $\mathcal{O}^{\prime}Hc$ an extension of an equivalence between $\mathcal{O}Gb$ and $\mathcal{O}Hc$? … [This] may be viewed as evidence for a refined version of the Abelian defect group conjecture, namely that for any $p$-modular system $(K,\mathcal{O},k)$ and any block $b$ of $\mathcal{O}G$ with abelian defect group $P$ and Brauer correspondent $c$, there is a splendid Rickard equivalence between $\mathcal{O}Gb$ and $\mathcal{O}N_{G}(H)c$.”
We do so by showing that Harris–Linckelmann’s construction in [HL00] can be adapted to a $p$-modular system with $k=\mathbb{F}_{p}$. One may use unique lifting of splendid Rickard equivalences to lift to $\mathcal{O}=\mathbb{Z}_{p}$, which then implies the result for any extension of the $p$-modular system $(\mathbb{Q}_{p},\mathbb{Z}_{p},\mathbb{F}_{p})$. In [HL00], the authors specify a $p$-modular system with $k$ algebraically closed. However, for many of the statements, one may replace the assumption that $k$ is algebraically closed with the assumption that $k$ is perfect with enough roots of unity, which we will use frequently.
For this section, we will refer to “blocks” as the primitive idempotents of $Z(\overline{\mathbb{F}}_{p}G)$ (note these always exist over some finite field), and refer to block idempotents over a finite field $k$ which may not be primitive upon extending scalars as “block idempotents of kGkG.” Note that for these block idempotents, defect groups are defined in the usual way (for instance, a maximal $p$-subgroup $P$ of $G$ for which $\operatorname{Br}_{P}(b)\neq 0$) and satisfy essentially all the usual properties. Brauer’s first main theorem asserts that there is a canonical bijection between the set of block idempotents of $kG$ with defect group $P$ and the set of block idempotents of $kN_{G}(P)$ with defect group $P$. Given a block idempotent $b$ of $kG$ with defect group $P$, the block idempotent $c$ of $kN_{G}(P)$ corresponding to $b$ under this bijection is called the Brauer correspondent.
@@
An $\mathcal{O}^{\times}$-group is a group $\hat{G}$ endowed with a group homomorphism $\mathcal{O}^{\times}\to\hat{G}$ whose image lies in Z(G^)Z($\hat${G}); we usually denote by $\hat{\lambda}$ the image of $\lambda\in\mathcal{O}^{\times}$ in $\hat{G}$, and by $\check{G}$ the opposite $\mathcal{O}^{\times}$-group, which as an abstract group is equal to $\hat{G}$ but endowed with the homomorphism $\mathcal{O}^{\times}\to\check{G}$ sending $\lambda\in\mathcal{O}^{\times}$ to $\check{\lambda}=(\hat{\lambda})^{-1}.$
@@
For our context, $\mathcal{O}^{\times}$-groups arise from the action of $G$ on a matrix algebra $S$ over $\mathcal{O}$ by algebra automorphisms. Denote by $^{x}s$ the action of $x\in G$ on $s\in S$. By the Skolem-Noether theorem, any automorphism of $S$ is inner, and hence for any $x\in G$, there is $s\in S$ such that $^{x}t=sts^{-1}$ for all $t\in S$. The group $\hat{G}$ of all such pairs $(x,s)\in G\times S^{\times}$ satisfying $^{x}t=sts^{-1}$ for all $t\in S$ becomes an $\mathcal{O}^{\times}$-group with the group homomorphism sending $\lambda\in\mathcal{O}^{\times}$ to $\hat{\lambda}=(1_{G},\lambda 1_{S})$. We call $\hat{G}$ the $\mathcal{O}^{\times}$-group defined by the action of $G$ in $S$. In fact, $\hat{G}$ is a central $\mathcal{O}^{\times}$-extension of $G$, as we have a short exact sequence
$1\to\mathcal{O}^{\times}\to\hat{G}\to G\to 1.$
@@
[L181, Corollary 5.3.4] Let $(K,\mathcal{O},k)$ be a $p$-modular system with $k$ algebraically closed. If $S$ has $\mathcal{O}$-rank prime to $p$ and $P$ is a finite $p$-group acting on a matrix algebra $S$, the action of $P$ lifts to a unique group homomorphism $\sigma\colon P\to S^{\times}$ such that $\det(\sigma(x))=1$ for all $x\in P$. In particular, $\hat{P}$ splits canonically as a direct product $\hat{P}\cong P\times\mathcal{O}^{\times}$, with the map sending $u\in P$ to $(u,\sigma(u))\in\hat{P}$ being a section of the canonical surjection $\hat{P}\to P$.
@@
For every $S$ and $P$ satisfying the hypotheses in the previous proposition, there exists a $p$-modular system $(K,\mathcal{O},k)$ with $k$ finite for which the previous proposition holds as well. See [L181, Remark 5.3.11] and [L181, Remark 1.2.19] for more details.
@@
[…]
Let $(K,\mathcal{O},k)\subseteq(K^{\prime},\mathcal{O}^{\prime},k^{\prime})$ be an extension of $p$-modular systems, let $S$ be a matrix algebra over $\mathcal{O}$, and set $S^{\prime}:=\mathcal{O}^{\prime}\otimes_{\mathcal{O}}S$. Let $\hat{G}_{\mathcal{O}}$ (resp. $\hat{G}_{\mathcal{O}^{\prime}}$) denote the $\mathcal{O}^{\times}$-group (resp. $(\mathcal{O}^{\prime})^{\times}$-group) defined by the action of $G$ on $S$ (resp. $G$ on $S^{\prime}$). Then, the obvious inclusion of $\mathcal{O}^{\prime}$-algebras $\mathcal{O}^{\prime}\otimes_{\mathcal{O}}\mathcal{O}_{*}\hat{G}_{\mathcal{O}}\to\mathcal{O}_{*}^{\prime}\hat{G}_{\mathcal{O}^{\prime}}$ is an isomorphism, since it sends the $\mathcal{O}^{\prime}$-basis given by representatives of $G$ in $\hat{G}$ to the analogous $\mathcal{O}^{\prime}$-basis.
@@
[…]
@@
To prove Broué’s conjecture holds, Harris–Linckelmann provide a reduction to the case where $G=O_{p^{\prime},p,p^{\prime}}(G)$ with blocks of principal type, then use structural properties of groups of this form to construct a splendid Rickard autoequivalence built from a Morita equivalence and a diagonally induced endosplit $p$-permutation resolution. We verify that their construction descends to the block’s field of realization. In [HL00], the construction is performed over $\mathcal{O}$, however, we will work in positive characteristic instead, then invoke Rickard’s lifting theorem, [R96, Theorem 5.2].
We first restate the structural theorem that describes block decompositions in the restricted case. In the original paper, this is stated over $\mathcal{O}$, but it is straightforward that this also holds over $k$.
@@
    The action of $P$ on $S$ lifts to a unique group homomorphism $\sigma\colon P\to S^{\times}$ such that $\det(\sigma(x))=1$ for all $x\in P$, and there is a $p^{\prime}$-subgroup $E$ of $\operatorname{Aut}(P)$ and a central $k^{\times}$-extension $\hat{E}$ of $E$ such that we have an isomorphism of $k^{\times}$-groups
@@
The assumption that $k$ is algebraically closed is invoked for Proposition 5.4, [HL00, Lemma 2.7] which is a consequence of Dade’s splitting theorem of fusion in endopermutation modules, and in the assumption that $S$ is a matrix algebra. One may verify Dade’s theorem satisfies the same large enough finite field replacement condition via a similar argument; see the proof of [L182, Theorem 7.9.1] and [L181, Remark 1.2.19]. Therefore, there exists a $p$-modular system $(K,\mathcal{O},k)$ with $k$ finite (large enough for $G$) for which the previous proposition holds as well.
@@
Therefore, we have a splendid Morita equivalence between $kHc$ and $k_{\ast}\hat{L}$, since source algebra equivalences induce splendid Morita equivalences (see [L01, Theorem 4.1], note that the notion of splendid-ness used in this article is stronger than Rickard’s original notion). Therefore, to verify Broué’s conjecture, it suffices to construct a splendid Rickard equivalence for $kGb$ and $k_{\ast}\hat{L}$. To verify the refined abelian defect group conjecture, it suffices to show that the equivalence descends to $\mathbb{F}_{p}[b]=\mathbb{F}_{p}[c]$ (see [BY22, Lemma 4.2]), then apply [KL18, Theorem 6.5(b)].
@@
    $D_{k}(P)=D^{\Omega}(P)$, where $D^{\Omega}(P)$ is the subgroup of $D_{k}(P)$ generated by equivalence classes of relative syzygies, that is, kernels of augmentation homomorphisms $kX\to k$, where $X$ is a transitive $P$-set.
@@
    If $M$ is an indecomposable endopermutation $kP$-module, then $M$ is absolutely indecomposable.
In particular, every endopermutation $kP$-module descends to $\mathbb{F}_{p}$ (or any intermediate field extension), that is, there exists an endopermutation $\mathbb{F}_{p}P$-module $\tilde{M}$ such that $k\otimes_{\mathbb{F}_{p}}\tilde{M}\cong M$.
@@
Statement (a) follows from the classification in [Bo06, Theorem 9.5] (clarifying the original classification in [D78a]), and statement (b) follows from [D78a, Theorem 6.6]. For the final statement, see [L182, Remark 7.8.5]. ∎
@@
[…]
@@
[…]
  • Since $V$ is an endopermutation $kP$-module with $P$ abelian, by [R96, Theorem 7.2] $V$ has an endosplit $p$-permutation resolution $X_{V}$. The uniqueness of the lifting $\sigma\colon P\to S^{\times}$ of the action of $P$ on $S$ with determinant 1 implies the isomorphism class of $V$ is $E$-stable, and thus $X_{V}$ is an $E$-stable endosplit $p$-permutation resolution. Then $V^{*}\otimes_{k}X_{V}$ is an endosplit (non-$p$-permutation) resolution of $V^{*}\otimes_{k}V$, and it follows from [HL00, Lemma 1.5(ii)] that $\operatorname{Ind}^{L}_{P}(V^{*}\otimes_{k}X_{V})$ is an endosplit resolution of $\operatorname{Ind}^{L}_{P}(V^{*}\otimes_{k}V),$ where $L:=G/O_{p^{\prime}}(G)$. Since $V$ is capped, $V^{*}\otimes_{k}V$ has the trivial $kP$-module as a direct summand, thus $\operatorname{Ind}^{L}_{P}(V^{*}\otimes_{k}V)$ has the trivial $kL$-module as a direct summand, since $E$ is a $p^{\prime}$-group. Therefore, $\operatorname{Ind}^{L}_{P}(V^{*}\otimes_{k}X_{V})$ has a direct summand $Y$ which is an endosplit resolution of the trivial $kL$-module. Setting
@@
[…]